# Nonequilibrium Probes of Quantum Geometry in Gapless Systems

### v1

@@ 5 Rotating drives in $\mathrm{SL}(2,\mathbb{R})$
The simplest inhomogeneous Hamiltonians (1.1) have a velocity profile involving a single wavenumber $k$, so that $v_{t}(x)=A_{t}+B_{t}\cos(2\pi kx/L-C_{t})$ with generally time-dependent coefficients $A_{t}>|B_{t}|$, $B_{t}$, and $C_{t}$. In contrast to the protocols of Secs. 3 and 4, such drives do not probe the full parameter space of conformal frames. They are limited, instead, to the subgroup of $\operatorname{Diff}(S^{1})$ that consists of transformations of the form
$e^{ikf(\varphi)}=\frac{\alpha e^{ik\varphi}+\beta}{\bar{\beta}e^{ik\varphi}+\bar{\alpha}}$ (5.1)
@@
where $\alpha,\beta\in\mathbb{C}$ satisfy $|\alpha|^{2}-|\beta|^{2}=1$. This subgroup is finite-dimensional and locally isomorphic to $\mathrm{SL}(2,\mathbb{R})\cong\mathrm{SU}(1,1)$. As a result, the dynamics of coherent states can be wholly expressed in terms of finite-dimensional matrices. We now study one such situation with a rotating drive $v_{t}(x)=A+B\cos(2\pi kx/L-\omega t)$ (where $A>|B|$ and $B$ are constants) and show that the evolution operator can be computed exactly for any driving frequency $\omega>0$ and general amplitudes $A$, $B$. This, in turn, allows us to compute nonlinear responses and return amplitudes in any driving regime, as illustrated in Fig. 3.
5.1 Evolution operator
@@
Pick an integer wavenumber $k$ and a real ‘rapidity’ parameter $\lambda>0$. Given the CFT stress-energy tensor ${\mathcal{T}}(\varphi)$, consider the time-dependent inhomogeneous Hamiltonian
@@
$H(t)=\frac{2\pi}{L}\oint\textrm{d}\varphi\,\Bigl[\cosh(2\lambda)-\sinh(2\lambda)\cos(k[\varphi-\omega t])\Bigr]{\mathcal{T}}(\varphi)+\frac{\pi c}{12L}[\cosh(2\lambda)-1].$ (5.2)
@@
This takes the form (2.15) for a rotating drive (4.13) whose ‘initial deformation’ $f_{0}$ is given by Eq. (5.1) with $\alpha=\cosh(\lambda)$ and $\beta=\sinh(\lambda)$. Equivalently, one has
@@
$H(t)=\frac{2\pi}{L}\biggl[\cosh(2\lambda)L_{0}-\frac{1}{2}\sinh(2\lambda)\biggl(e^{ik\omega t}L_{k}+e^{-ik\omega t}L_{-k}\biggr)\biggr]+\frac{\pi c}{12L}\bigl[(k^{2}-1)\cosh(2\lambda)-1\bigr]$ (5.3)
@@
in terms of the Virasoro generators introduced in Eq. (2.7). The Hamiltonian thus involves only $\{L_{-k},L_{0},L_{k}\}$, which generate an $\mathfrak{sl}(2,\mathbb{R})\cong\mathfrak{su}(1,1)$ subalgebra (see Appendix D). It follows that the evolution operator
$U(t)\equiv\overset{\longleftarrow}{{\cal T}}\exp\left(-i\int_{0}^{t}\textrm{d}t^{\prime}\,H(t^{\prime})\right),$ (5.4)
@@
with ordering $\overset{\longleftarrow}{{\cal T}}$ so that time increases from right to left, belongs to the corresponding $\mathrm{SL}(2,\mathbb{R})\cong\mathrm{SU}(1,1)$ subgroup of $\operatorname{Diff}(S^{1})$. Since this is a matrix group, one may hope to use matrix multiplication to compute $U(t)$, despite the nontrivial time ordering $\overset{\longleftarrow}{{\cal T}}$. No such simplification would be available in the full Virasoro algebra.
To obtain the evolution operator (5.4), we write the time-ordered exponential as
$U(t)=\lim_{M\to\infty}\prod_{m=0,\ldots,M}^{\longleftarrow}e^{-iH(m\delta t)\delta t},$ (5.5)
@@
$U(t)=e^{-i\frac{\pi c}{12L}k^{2}t[\cosh(2\lambda)-1]}e^{-i\omega t(L_{0}-c/24)}e^{-i\frac{2\pi}{L}t\bigl([\cosh(2\lambda)-L/T](L_{0}-c/24)-\sinh(2\lambda)\frac{L_{-k}+L_{k}}{2}\bigr)},$ (5.6)
expressed in terms of the Virasoro generators. Note the structure of this expression, which is consistent with the evolution operator for coherent states in Eq. (4.4) and the specific form (4.14) of $g_{t}$ for rotating drives. Indeed, using $e^{i\alpha_{t}}\,{\cal U}(g_{t})=U(T)\,{\cal U}(g_{0})$, one can identify the second exponential in Eq. (5.6) with the undeformed rotation $R_{\omega t}$, while the last exponential in Eq. (5.6) can be shown to represent the ‘deformed rotation’ $\zeta\circ R_{\Omega t}\circ\zeta^{-1}$. In the case at hand, $\zeta$ is an $\mathrm{SU}(1,1)$ deformation of the form (5.1) with (see again Appendix D)
$\alpha=\frac{1+\Lambda}{2\sqrt{\Lambda}},\qquad\beta=\frac{1-\Lambda}{2\sqrt{\Lambda}},\qquad\Lambda\equiv 2\pi\frac{e^{-2\lambda}-L/T}{L\Omega},$ (5.7)
@@
$\Omega=\frac{2\pi}{L}\sqrt{1-2\cosh(2\lambda)(L/T)+(L/T)^{2}},$ (5.8)
all of which is consistent with Eq. (4.15). What is special about $\mathrm{SL}(2,\mathbb{R})$ drives is that the deformed rotation can be written as an explicit exponential of finitely many Virasoro generators. This would not be true with more general rotating drives, as we saw in Sec. 4.
@@
Since the solution (4.4) of the Schrödinger equation consists of coherent states, the dynamics encoded by the evolution operator (5.6) defines a path in the parameter space (2.20): at time $t$, the state vector $|\psi(t)\rangle=e^{i\alpha_{t}}\,{\cal U}(g_{t})|h\rangle$ defines a unique coset $g_{t}G_{0}\in G/G_{0}$. (Conversely, the coset $g_{t}G_{0}$ uniquely determines the ray $|\psi(t)]=[\,{\cal U}(g_{t})|h\rangle]$.) The situation is especially simple for $\mathrm{SL}(2,\mathbb{R})$ drives, which only involve three Virasoro generators $\{L_{-k},L_{0},L_{k}\}$ and therefore probe a two-dimensional parameter space $\mathrm{SL}(2,\mathbb{R})/S^{1}$ (i.e. a hyperbolic plane). One can thus view the motion of the ray $|\psi(t)]$ as a curve on a hyperbolic disk, as in Fig. 5. Concretely, the curve is obtained by writing the ray of the state vector as
$\bigl[|\psi(t)\rangle\bigr]=\bigl[e^{\chi_{t}L_{-k}-\bar{\chi}_{t}L_{k}}|h\rangle\bigr]$ (5.9)
@@
where $\chi_{t}$ is a time-dependent complex number that can be conveniently parametrized as $\chi_{t}\equiv\frac{1}{k}\operatorname{artanh}(|\eta_{t}|)\frac{\eta_{t}}{|\eta_{t}|}$ in terms of $\eta_{t}$ on the unit (Poincaré) disk. One finds (see Appendix D)
$\eta_{t}\equiv\tanh(k|\chi_{t}|)\frac{\chi_{t}}{|\chi_{t}|}=\tanh(\lambda)\frac{1+i\Bigl(\frac{2\pi}{L\Omega}+\frac{\omega}{\Omega}\Bigr)\tan(k\Omega t/2)}{1+i\Bigl(\frac{2\pi}{L\Omega}-\frac{\omega}{\Omega}\Bigr)\tan(k\Omega t/2)}e^{-ik\omega t}$ (5.10)
@@
with $\Omega$ the frequency in Eq. (5.8). This is illustrated in Fig. 5. Note the behavior in the adiabatic limit, where $\lim_{L/T\to 0}\eta_{t}=\tanh(\lambda)e^{-ik\omega t}$ essentially coincides with the rotating drive, as was to be expected. In fact, the complex variable $\eta_{t}$ is known to satisfy a Riccati equation with time-dependent coefficients determined by the external drive [63], but this perspective is not needed in our case.
@@
Figure 5: Time evolution of the unit disk coordinate $\eta_{t}$ in Eq. (5.10) induced by the dynamics given by the Hamiltonian (5.3) with $k=1$ and $\lambda=0.2$. The result $\tanh(\lambda)e^{-ik\omega t}$ in the adiabatic limit $L/T\to 0$ for the trajectory $\eta_{t}$ is shown as a black dashed line, while the blue curve shows the nonadiabatic result for $L/T=1/15$. After one full period, the final state (red dot) does not exactly return to the initial state (blue dot). The slight mismatch between the corresponding wave functions (coherent states) gives rise to a nearly saturated return probability.
5.2 Nonlinear and linear responses
@@
A first application of the evolution operator (5.6) is to compute the effect of the driven Hamiltonian (5.2) on any operator (2.6) generated by the stress-energy tensor, similarly to what we did in Sec. 3 to linear order. The result is actually quite general: provided $g_{t}$ is chosen such that the state vector $|\phi(t)\rangle=e^{i\alpha_{t}}\,{\cal U}(g_{t})|h\rangle$ solves the Schrödinger equation (4.3) with initial condition $|\phi(0)\rangle=\,{\cal U}(g_{0})|h\rangle$, the corresponding expectation value of $i\mathfrak{u}(\mathbf{Y})$ is
@@
$\langle\phi(t)|i\mathfrak{u}(\mathbf{Y})|\phi(t)\rangle=\frac{1}{2\pi}\oint\textrm{d}\varphi\,Y(\varphi)\biggl(\Bigl(h-\frac{c}{24}\Bigr)\bigl[(g_{t}^{-1})^{\prime}(\varphi)\bigr]^{2}-\frac{c}{12}\{g_{t}^{-1}(\varphi),\varphi\}\biggr).$ (5.11)
@@
This readily follows from the transformation law (2.14) of the CFT stress tensor, and it holds regardless of the initial $g_{0}$. In the case of rotating drives, Eq. (4.14) allows one to express $\langle\phi(t)|i\mathfrak{u}(\mathbf{Y})|\phi(t)\rangle$ in terms of the deformation $\zeta(\varphi)$, the frequency $\Omega$, and $g_{0}$.
@@
For the specific case of rotating drives in $\mathrm{SL}(2,\mathbb{R})$, this gives explicit access to the full nonlinear response. Indeed, $\zeta(\varphi)$ is explicitly known to be of the form (5.1) with $\alpha$, $\beta$ in Eq. (5.7), and $\Omega$ is also known exactly from Eq. (5.8). Choosing $g_{0}$ to be the identity, and considering the special case of small perturbations, i.e. when $\epsilon=\lambda$ is small, it follows that
$\langle\phi(t)|{\cal O}|\phi(t)\rangle=\Bigl(h-\frac{c}{24}\Bigr)\hat{Y}_{0}-i\epsilon{\cal F}_{k,-k}\left(\hat{Y}_{k}\frac{e^{ik\omega t}-e^{ik2\pi t/L}}{k(1-L/T)}+\hat{Y}_{-k}\frac{e^{-ik\omega t}-e^{-ik2\pi t/L}}{k(1-L/T)}\right)+O(\epsilon^{2})$ (5.12)
@@
in terms of the Berry curvature ${\cal F}_{m,n}$ of Eq. (2.23b). This agrees with the general linear-response result (3.8)–(3.9) by noting that the $\mathbf{X}_{t}$ in Eq. (3.2) is now given by $X_{t}(\varphi)=-\frac{2}{k}\sin(k[\varphi-\omega t])$. In fact, the general expression (5.11) is consistent with our linear-response formulas (3.8)–(3.9): this can be shown by expressing $g_{t}(\varphi)$ in terms of a weakly perturbed $f_{t}(\varphi)=\varphi+\epsilon X_{t}(\varphi)$.
5.3 Return amplitudes and probabilities

### v2

@@ V Rotating drives in $\mathrm{\bf SL}\bm{(}\bm{2}\bm{,}\mathbb{R}\bm{)}$
The simplest case of an inhomogeneous Hamiltonian (1) has a velocity profile involving a single wavenumber $k$, so that $v_{t}(x)=A_{t}+B_{t}\cos(2\pi kx/L-C_{t})$ with generally time-dependent coefficients $A_{t}>|B_{t}|$, $B_{t}$, and $C_{t}$. In contrast to the protocols of Secs. III and IV, such drives do not probe the full parameter space of conformal frames. They are limited, instead, to the subgroup of $\operatorname{Diff}(S^{1})$ that consists of transformations of the form
$e^{ikf(\varphi)}=\frac{\alpha e^{ik\varphi}+\beta}{\bar{\beta}e^{ik\varphi}+\bar{\alpha}}$ (71)
@@
for $\alpha,\beta\in\mathbb{C}$ satisfying $|\alpha|^{2}-|\beta|^{2}=1$. This subgroup is finite-dimensional and locally isomorphic to $\mathrm{SL}(2,\mathbb{R})\cong\mathrm{SU}(1,1)$. As a result, the dynamics of coherent states can be wholly expressed in terms of finite-dimensional matrices. We now study one such situation with a rotating drive $v_{t}(x)=A+B\cos(2\pi kx/L-\omega t)$ (where $A>|B|$ and $B$ are constants) and show that the evolution operator can be computed exactly for any driving frequency $\omega>0$ and general amplitudes $A$, $B$. This, in turn, allows us to compute nonlinear responses and return amplitudes in any driving regime, as illustrated in Fig. 3.
V.1 Evolution operator
@@
Pick an integer wavenumber $k$ and a real ‘rapidity’ parameter $\lambda>0$. Given the component ${\mathcal{T}}(\varphi)$ of the CFT stress-energy tensor, consider the time-dependent inhomogeneous Hamiltonian
@@
$H(t)=\frac{2\pi}{L}\oint\textrm{d}\varphi\,\Bigl[\cosh(2\lambda)-\sinh(2\lambda)\cos(k[\varphi-\omega t])\Bigr]{\mathcal{T}}(\varphi)+\frac{\pi c}{12L}\bigl[\cosh(2\lambda)-1\bigr].$ (72)
@@
This takes the form (21) for a rotating drive (55) whose ‘initial deformation’ $f_{0}$ is given by Eq. (71) with $\alpha=\cosh(\lambda)$ and $\beta=\sinh(\lambda)$. Equivalently,
@@
$H(t)=\frac{2\pi}{L}\biggl[\cosh(2\lambda)L_{0}-\frac{1}{2}\sinh(2\lambda)\Bigl(e^{ik\omega t}L_{k}+e^{-ik\omega t}L_{-k}\Bigr)\biggr]+\frac{\pi c}{12L}\bigl[(k^{2}-1)\cosh(2\lambda)-1\bigr]$ (73)
@@
using the Virasoro generators $L_{n}$ in Eqs. (13)–(15). The Hamiltonian thus involves only $\{L_{-k},L_{0},L_{k}\}$, which generate an $\mathfrak{sl}(2,\mathbb{R})\cong\mathfrak{su}(1,1)$ subalgebra (see Appendix D). It follows that the evolution operator
$U(t)\equiv\overset{\longleftarrow}{{\cal T}}\exp\left(-i\int_{0}^{t}\textrm{d}t^{\prime}\,H(t^{\prime})\right),$ (74)
@@
with ordering $\overset{\longleftarrow}{{\cal T}}$ so that time increases from right to left, belongs to the corresponding $\mathrm{SL}(2,\mathbb{R})\cong\mathrm{SU}(1,1)$ subgroup of $\operatorname{Diff}(S^{1})$. Since this is a matrix group, one may hope to use matrix multiplication to compute $U(t)$, despite the nontrivial time ordering. No such simplification would be available in the full Virasoro algebra.
To obtain the evolution operator (74), write the time-ordered exponential as
$U(t)=\lim_{M\to\infty}\prod_{m=0,\ldots,M}^{\longleftarrow}e^{-iH(m\delta t)\delta t},$ (75)
@@
$U(t)=e^{-i\frac{\pi c}{12L}k^{2}t[\cosh(2\lambda)-1]}e^{-i\omega t(L_{0}-c/24)}$ (76)
$\quad\times e^{-i\frac{2\pi}{L}t\bigl([\cosh(2\lambda)-L/T](L_{0}-c/24)-\sinh(2\lambda)\frac{L_{-k}+L_{k}}{2}\bigr)},$
expressed in terms of the Virasoro generators. Note the structure of this expression, which is consistent with the evolution operator for coherent states in Eq. (47) and the specific form (56) of $g_{t}$ for rotating drives. Indeed, using $e^{i\alpha_{t}}\,{\cal U}(g_{t})=U(T)\,{\cal U}(g_{0})$, one can identify the second exponential in Eq. (76) with the undeformed rotation $R_{\omega t}$, while the last exponential in Eq. (76) can be shown to represent the ‘deformed rotation’ $\zeta\circ R_{\Omega t}\circ\zeta^{-1}$. In the case at hand, $\zeta$ is an $\mathrm{SU}(1,1)$ deformation of the form (71) with (see again Appendix D)
$\alpha=\frac{1+\Lambda}{2\sqrt{\Lambda}},\quad\;\beta=\frac{1-\Lambda}{2\sqrt{\Lambda}},\quad\;\Lambda\equiv 2\pi\frac{e^{-2\lambda}-L/T}{L\Omega},$ (77)
@@
$\Omega=\frac{2\pi}{L}\sqrt{1-2\cosh(2\lambda)(L/T)+(L/T)^{2}},$ (78)
all of which is consistent with Eq. (57). What is special about $\mathrm{SL}(2,\mathbb{R})$ drives is that the deformed rotation can be written as an explicit exponential of finitely many Virasoro generators. This would not be true with more general rotating drives, as we saw in Sec. IV.
@@
Since the solution (47) of the Schrödinger equation consists of coherent states, the dynamics encoded by the evolution operator (76) defines a path in the parameter space (26): at time $t$, the state vector $|\psi(t)\rangle=e^{i\alpha_{t}}\,{\cal U}(g_{t})|h\rangle$ defines a unique coset $g_{t}G_{0}\in G/G_{0}$. (Conversely, the coset $g_{t}G_{0}$ uniquely determines the ray $|\psi(t)]=[\,{\cal U}(g_{t})|h\rangle]$.) The situation is especially simple for $\mathrm{SL}(2,\mathbb{R})$ drives, which only involve three Virasoro generators $\{L_{-k},L_{0},L_{k}\}$ and therefore probe a two-dimensional parameter space $\mathrm{SL}(2,\mathbb{R})/S^{1}$ (i.e. a hyperbolic plane). One can thus view the motion of the ray $|\psi(t)\rangle]$ as a curve on a hyperbolic disk, as in Fig. 5. Concretely, the curve is obtained by writing the ray of the state vector as
$\bigl[|\psi(t)\rangle\bigr]=\bigl[e^{\chi_{t}L_{-k}-\bar{\chi}_{t}L_{k}}|h\rangle\bigr]$ (79)
@@
for $\chi_{t}$ a time-dependent complex function that can be conveniently parametrized as $\chi_{t}\equiv\frac{1}{k}\operatorname{artanh}(|\eta_{t}|)\frac{\eta_{t}}{|\eta_{t}|}$ in terms of $\eta_{t}$ on the unit (Poincaré) disk. One finds (see Appendix D)
$\eta_{t}$ $\equiv\tanh(k|\chi_{t}|)\frac{\chi_{t}}{|\chi_{t}|}$ (80)
$=\tanh(\lambda)\,\frac{1+i\Bigl(\frac{2\pi}{L\Omega}+\frac{\omega}{\Omega}\Bigr)\tan(k\Omega t/2)}{1+i\Bigl(\frac{2\pi}{L\Omega}-\frac{\omega}{\Omega}\Bigr)\tan(k\Omega t/2)}\,e^{-ik\omega t}$
@@
with $\Omega$ the frequency (78). This is illustrated in Fig. 5. Note the behavior in the adiabatic limit, where $\lim_{L/T\to 0}\eta_{t}=\tanh(\lambda)e^{-ik\omega t}$ essentially coincides with the rotating drive, as was to be expected. In fact, the complex variable $\eta_{t}$ is known to satisfy a Riccati equation with time-dependent coefficients determined by the external drive [94], but this perspective is not needed in our case.
@@
Figure 5: Time evolution of the unit disk coordinate $\eta_{t}$ in Eq. (80) induced by the dynamics given by the Hamiltonian (73) with $k=1$ and $\lambda=0.22$. The result $\tanh(\lambda)e^{-ik\omega t}$ in the adiabatic limit $L/T\to 0$ for the trajectory $\eta_{t}$ is shown as a black dashed line, while the blue curve shows the nonadiabatic result for $L/T=1/11$. After one full period, the final state (red dot) does not exactly return to the initial state (blue dot). The slight mismatch between the corresponding coherent states gives rise to a nearly saturated return probability.
V.2 Nonlinear and linear responses
@@
A first application of the evolution operator (76) is to compute the effect of the driven Hamiltonian (72) on any operator (12) generated by the stress-energy tensor, similarly to what we did in Sec. III to linear order. The result is actually quite general: provided $g_{t}$ is chosen such that the state vector $|\phi(t)\rangle=e^{i\alpha_{t}}\,{\cal U}(g_{t})|h\rangle$ solves the Schrödinger equation (46) with initial condition $|\phi(0)\rangle=\,{\cal U}(g_{0})|h\rangle$, the corresponding expectation value of $i\mathfrak{u}(\mathbf{W})$ is
@@
$\langle\phi(t)|i\mathfrak{u}(\mathbf{W})|\phi(t)\rangle=\frac{1}{2\pi}\oint\textrm{d}\varphi\,W(\varphi)$ (81)
$\quad\times\biggl(\Bigl(h-\frac{c}{24}\Bigr)\bigl[(g_{t}^{-1})^{\prime}(\varphi)\bigr]^{2}-\frac{c}{12}\{g_{t}^{-1},\varphi\}\biggr).$
@@
This follows from the transformation law (20) of the CFT stress tensor, and it holds regardless of the initial $g_{0}$. In the case of rotating drives, Eq. (56) allows one to express $\langle\phi(t)|i\mathfrak{u}(\mathbf{W})|\phi(t)\rangle$ in terms of the deformation $\zeta(\varphi)$, the frequency $\Omega$, and $g_{0}$.
@@
For the specific case of rotating drives in $\mathrm{SL}(2,\mathbb{R})$, this gives explicit access to the full nonlinear response. Indeed, $\zeta(\varphi)$ is known to be of the form (71) with $\alpha$, $\beta$ in Eq. (77), and $\Omega$ is also known exactly from Eq. (78). Choosing $g_{0}$ to be the identity, and considering the special case of small perturbations $\epsilon=\lambda\ll 1$, it follows that
$\langle\phi(t)|i\mathfrak{u}(\mathbf{W})|\phi(t)\rangle=\Bigl(h-\frac{c}{24}\Bigr)\hat{W}_{0}$ (82)
$\quad\begin{aligned} &-i\epsilon{\cal F}_{k,-k}\biggl(\hat{W}_{k}\frac{e^{ik\omega t}-e^{ik2\pi t/L}}{k(1-L/T)}\\ &\qquad\qquad\quad+\hat{W}_{-k}\frac{e^{-ik\omega t}-e^{-ik2\pi t/L}}{k(1-L/T)}\biggr)+O(\epsilon^{2})\end{aligned}$
@@
in terms of the Berry curvature ${\cal F}_{m,n}$ of Eq. (30b). This agrees with the general linear-response result (39)–(40) by noting that the $\mathbf{X}_{t}$ in Eq. (33) is now given by $X_{t}(\varphi)=-\frac{2}{k}\sin(k[\varphi-\omega t])$. In fact, the general expression (81) is consistent with our linear-response formulas (39)–(40): this can be shown by expressing $g_{t}(\varphi)$ in terms of a weakly perturbed $f_{t}(\varphi)=\varphi+\epsilon X_{t}(\varphi)$.
V.3 Return amplitudes and probabilities